Sprout:

I don't want the answer, I want to know how to solve this. (Give me a minute to type it out.)

Sprout:

\[x ^{4}\times{x}{\times}x ^{3}\]

Extrinix:

\(x^a\times x^b\times x^b = x^{\space a+\space b\space+\space c}\)

Extrinix:

That's the exponent rule for multiplication.

Sprout:

so 4+1+3?

Extrinix:

Correct

Sprout:

So the answer is \[x ^{8}\]?

Extrinix:

Correct

Sprout:

Thanks. :3

Extrinix:

Brief overview of the rules themselves: Multiplication: \(x^a \times x^b = x^{a+b}\) Division: \(\dfrac{x^a}{x^b} = x^{a-b}\)